what is the system of simultaneous equations by using the substitution method?
x=10-y
y=x-8

Answer by sofiyac(983)

what is the system of simultaneous equations by using the substitution method?
x=10-y
y=x-8
I think you meant to say "solve the system of equations using the substitution method".
Ok so you already have that y=x-8 in the second equation, so plug that into your first one, you get
x=10-(x-8) then distribute the minus and you get
x=10-x+8 simplify and you get
x=18-x now add x to each side of the equation and you get
2x=18 divide each side by 2 and you get
x=9 then you plug that into the second original equation y=9-8 and you get
y=1
so the solution is (9,1)